# SEREP: Semantic Facial Expression Representation for Robust In-the-Wild Capture and Retargeting

### v2

@@ 4 Experiments on expression representation
The experiments in this section evaluate the semantic face expression model (Sec. 3.1). We detail its implementation and assess its performance on reconstructing expressions and preserving identity in retargeting.
@@ 4.1 Implementation detail
We train the expression model using two datasets: static neutral meshes, and expressive sequences. We use neutral scans from 865 individuals, and dynamic sequences from 10 subjects, each captured over sessions lasting 3 to 19 minutes. These sequences cover a variety of expressions and talking sequences. We evaluate our results on static scans of 56 subjects, including a neutral mesh and 19 expressions. All meshes share the same topology, with 13k vertices.
@@
In all experiments, we use the official implementation of each baseline taken from their respective code repository. For quantitative experiments, results reported in bold are statistically significant based on a Wilcoxon signed rank test ($\rho<0.001$).
@@ 4.2 Expression reconstruction
First, we evaluate the ability of the proposed method to reconstruct expressive face meshes from 56 identities from the test set. For each expressive mesh, we define an optimization problem: given the subject’s neutral and expressive meshes, find the expression coefficients that minimize the reconstruction loss. We compare our method to the widely used FLAME model [33], evaluating all 1064 expressive meshes from the test set.
@@ 4.3 Identity preservation
We now evaluate the identity preservation of the models on an expression retargeting task. For this evaluation, we consider all pairs of [source, target] subjects from the test set. For each expressive mesh, we obtain the optimal expression code (as in Sec. 4.2) and apply it to the target’s neutral mesh. In total, we obtain 61,600 retargeted meshes for each method. We render the retargeted meshes using the GT neutral texture and a 50mm camera using Blender. Finally, we measure the similarity between the rendered images of the GT and the retargeted images using the cosine similarity (CSIM) of identity embedding as commonly done in face re-enactment research [38, 15, 26]. A higher CSIM indicates a better preservation of the identity. In other words, a lower source identity traits leakage on the retargeted mesh.
@@ 5.1 MultiREX Benchmark
Inspired by the REALY benchmark [7], we adopt a region-based evaluation method, dividing the face into four regions and performing region-based rigid alignment before assessment. This avoids penalizing a model due to rigid misalignment between the predicted and GT meshes, and instead focuses on the non-rigid deformations. The regions we evaluate are the forehead, cheek, mouth, and nose. For each region, we find the optimal rigid alignment between the GT and predicted meshes in the Multiface topology and compute the per-vertex error. Masks and alignment details are provided in the supplementary material.
We publicly release [^1] Link to code and assets - Coming soon: (i) the code to download assets, (ii) neutral meshes in the FLAME topology alongside the code to convert between FLAME and Multiface topologies, (iii) code to run the benchmark and compute the metrics.
@@ 5.2 Implementation details of the capture model
Method Cheek Forehead Mouth Nose Average
DECA [18] $3.54\pm 1.59$ $1.82\pm 0.70$ $3.71\pm 1.81$ $1.30\pm 0.47$ $2.59$
SMIRK [39] $4.13\pm 1.59$ $1.84\pm 0.66$ $3.69\pm 1.47$ $1.29\pm 0.43$ $2.74$
Ours $\mathbf{2.83\pm 1.28}$ $1.89\pm 0.81$ $\mathbf{2.95\pm 1.51}$ $\mathbf{1.08\pm 0.45}$ $2.19$
@@
In addition, we generate 135k synthetic images following the procedure described in Sec. 3.2. We project mesh vertices to camera coordinates to obtain ground truth landmarks in image space, except for eyebrows. Recent work [19] notes that eyebrow landmarks are not precise when obtained from mesh vertices of synthetic data. The location of eyebrows in a synthetic image depends on both mesh and texture information. In our work, we simply obtain eyebrow landmarks using a landmark detector [56].
@@ 5.3 Results on MultiREX
We now show quantitative results on geometric expression capture, on the proposed MultiREX benchmark. In all tables, results reported in bold are statistically significant based on a Wilcoxon signed rank test ($\rho<0.001$). We process videos from all subjects and camera views, comparing our model to DECA [18], EMICA [10, 11] and SMIRK [39]. For this experiment, we use ground truth bounding boxes, computed from projecting keypoint vertices to the image space.
For all models, we consider that the neutral mesh is given, therefore only accounting for geometric deformations due to the facial expressions. For a given source image $X$, with identity mesh $N_{s}$, we obtain the expressive mesh $M=\mathbf{D}_{\text{mesh}}(\mathbf{E}_{\text{exp}}(X),\mathbf{E}_{\text{id}}(N_{s}))$. For the other methods, we use their image encoders to obtain the expression coefficients $\boldsymbol{\Psi}$, and apply them directly to the neutral mesh $M=N_{t}+\boldsymbol{\Psi}\boldsymbol{E}$. For the FLAME-based models, this is followed by the Linear Blend Skinning (LBS) function for jaw rotation.
@@
We further analyze the viewpoint robustness of each method by reporting the average reconstruction errors for each viewing angle independently in Tab. 3. These results indicate that our model maintains performance better than competing methods as the viewing angle shifts from the frontal view. In contrast, the performance of EMICA and SMIRK decreases faster as the camera angle increases. This can be visualized in Fig. 5, where we show results for the same expression reconstructed from five different views. Additional examples are provided in the supplementary material. Our model produces consistent results across views, particularly in side views when compared to competing methods. This is especially apparent in the mouth shape of the subject. This is explained by the fact that our method relies on a semantic space that represents facial expressions in their entirety, rather than relying solely on image space information where some regions may be occluded.
5.4 In-the-wild retargeting results
@@
In this section, we compare our method with recent state-of-the-art techniques, DECA [17], EMICA [11], SMIRK [39], and 3DDFA-V3 [45], on the expression retargeting task from in-the-wild images. Since ground truth data for corresponding expressions across subjects on in-the-wild images is not possible to obtain, we compare methods qualitatively. We show several examples in Fig. 6, where the expressions from the images on the first column are transferred to the target identities of the second one.
This is done similarly to the procedure in Sec. 5.3, but here we consider a neutral mesh from another subject, $N_{t}$, obtained with an off-the-shelf reconstruction method [14]. 3DDFAv3 uses the Basel 3DMM, which does not use LBS for jaw movements, so this step is not applied.
@@ 6 Limitations
We focus on facial expression capture and consequently assume a known identity. We acknowledge that the methods used for comparison in our experiments, while being the best candidates to our knowledge, were designed to address the broader task of reconstructing facial shape, expression and textures.
@@ 7 Conclusion
In this work, we present SEREP, a 3D facial expression capture model that operates at a semantic level, enabling precise capture and retargeting of in-the-wild facial expressions. We introduce a semi-supervised learning framework in which we first train a semantic expression encoder using unpaired data, subsequently used to guide the training of our expression capture model. This capture model leverages ground truth expression codes used to generate synthetic image data, as well as facial landmarks from both synthetic and in-the-wild images, to enhance accuracy.
Our experiments underscore the strengths of our expression representation, demonstrating its effectiveness for precise expression capture and retargeting from in-the-wild images. We show that SEREP handles complex expressions and remains robust across various capture conditions, including diverse viewpoints. Additionally, we introduce MultiREX, a new benchmark that evaluates expression capture directly in 3D space, eliminating the measurement imprecision caused by projecting geometry back into image space, as done in previous evaluation protocols. For future work, we aim to make our capture model temporally aware to improve video capture quality and plan to integrate camera position information for enhanced supervision.
@@ 8 Acknowledgements
We would like to thank authors of the Multiface dataset [51], as well as authors of DECA [18], EMICA [11], SMIRK [39], and 3DDFAv3 [45] for releasing their code and advancing research in the field.
Appendix A Appendices
A.1 Implementation details
For training the semantic face expression model, we use the following importance weight in eq. 1 of the main manuscript. $\lambda_{rec}=0.001$, $\lambda_{cycle}=1.0$, $\lambda_{delta}=0.01$, $\lambda_{edge}=10000$ and $\lambda_{eyes}=0.01$. Details for the model’s architecture are provided Tab. 4.
For the Face capture model, the following importance weight are used in eq. 2 of the main manuscript: $\lambda_{code}=10$, $\lambda_{lmks}=1$ and $\lambda_{domain}=0.005$. For the gradient reversal, we use a scale factor equal to 1. Details for the model’s architecture are provided Tab. 5.
A.2 MultiREX additional details
@@
A.3 More comparisons
@@
A.3.1 Comparison on MultiREX benchmark
@@
A.3.2 Comparison on in-the-wild images
Here, we show more qualitative results for in-the-wild retargeting to other subjects and considering different source expressions. Our model is more faithful to the source expression in most scenarios while accounting for the face morphology and through its semantic expression model.
A.4 Synthetic dataset
Fig. 13 shows samples from our synthetic dataset used to train our face capture model. Rendering is done using Blender, with the Cycles renderer. We use randomly selected environment maps [^4] https://polyhaven.com. We use the same meshes and textures for teeth and eyes for all subjects. They are placed procedurally based on the vertex positions of the eyelids and jaw.

### v3

@@ 4.1 Implementation detail
We train the expression model using two datasets: static neutral meshes, and expressive sequences. We use neutral scans from 865 individuals, and dynamic sequences from 10 subjects, each captured over sessions lasting 3 to 19 minutes. These sequences cover a variety of expressions and talking sequences, which should reasonably cover the expression space naturally bounded by human physiology. We evaluate our results on static scans of 56 subjects, including a neutral mesh and 19 fixed expressions. All meshes share the same topology, with 13k vertices.
@@
In all experiments, we use the official implementation of each baselineFor quantitative experiments, results reported in bold are statistically significant based on a Wilcoxon signed rank test ($\rho<0.001$).
@@ 4.2 Expression reconstruction
First, we evaluate the ability of the proposed method to reconstruct expressive face meshes for the 56 identities of the test set (i.e., evaluating for 1064 meshes). For each expressive mesh, we define an optimization problem: given the subject’s neutral and expressive meshes, find the expression coefficients that minimize the reconstruction loss. We compare our method to the widely used FLAME model [31].
@@ 4.3 Identity preservation
For this evaluation, we consider all pairs of [source, target] subjects from the test set. For each expressive mesh, we optimize the expression code (as in Sec. 4.2) and apply it to the target’s neutral mesh. In total, we obtain 61,600 retargeted meshes for each method. We render the retargeted meshes using the GT neutral texture and a 50mm camera using Blender. Finally, we measure the similarity between the rendered images of the GT and the retargeted images using the cosine similarity (CSIM) of identity embedding as commonly done in face re-enactment research [35, 13, 25]. A higher CSIM indicates a better preservation of the identity. In other words, a lower source identity traits leakage on the retargeted mesh.
@@ 4.4 Ablations on the semantic expression model
We investigate removing either the delta regularization term (no delta reg.) or the identity conditioning from the decoder (no ID in decoder) to validate its impact on identity-expression separation. In the latter case, the model learns to generate vertex displacements solely from expression codes that are directly added to the neutral mesh, as commonly done in 3DMMs. We also investigate removing the edge preservation and eye closure loss in supplementary material. Those two mostly enforce visual quality. As shown in Fig. 5, without delta loss, the decoder collapses to mostly generating neutral faces, severely limiting expression reconstruction accuracy and causing significant identity leakage during retargeting (Tab. 1). While the no ID in decoder model achieves comparable reconstruction error, it learns additive displacements similar to delta transfer (Fig. 5). This compromise identity preservation in retargeting, and highlights the need for identity conditioning.
Figure 5: Ablations on the semantic expression model.
Method Cheek Forehead Mouth Nose Average
DECA [16] $3.54\pm 1.59$ $1.82\pm 0.70$ $3.71\pm 1.81$ $1.30\pm 0.47$ $2.59$
SMIRK [36] $4.13\pm 1.59$ $1.84\pm 0.66$ $3.69\pm 1.47$ $1.29\pm 0.43$ $2.74$
Ours (no $L_{\text{d}}$) $\mathbf{2.80\pm 1.33}$ $\mathbf{1.64\pm 0.66}$ $\mathbf{2.95\pm 1.59}$ $\mathbf{1.00\pm 0.42}$ $2.10$
Ours $2.83\pm 1.28$ $1.89\pm 0.81$ $\mathbf{2.95\pm 1.51}$ $1.08\pm 0.45$ $2.19$
@@ 5.1 MultiREX Benchmark
Inspired by the REALY benchmark [5], we adopt a region-based evaluation method, dividing the face into four regions (i.e., forehead, cheek, mouth, and nose) and performing region-based rigid alignment before assessment. This avoids penalizing a model due to rigid misalignment between the predicted and GT meshes, and instead focuses on the non-rigid deformations. For each region, we find the optimal rigid alignment between the GT and predicted meshes in the Multiface topology and compute the per-vertex error. Masks and alignment details are provided in the supplementary material.
We publicly release [^1] Link to code and assets: https://github.com/ubisoft/ubisoft-laforge-multirex: (i) the code to download assets, (ii) neutral meshes in the FLAME topology alongside the code to convert between FLAME and Multiface topologies, (iii) code to run the benchmark and compute the metrics.
@@ 5.2 Implementation details of the capture model
In addition, we generate 135k synthetic images with corresponding codes, following the procedure described in Sec. 3.2. We also project mesh vertices to camera coordinates to obtain ground truth landmarks in image space, except for eyebrows. Eyebrow landmarks are not precise for mesh vertices [17] as they also depend on texture for synthetic data, so we use a landmark detector to obtain them [52].
For the following experiments, given an identity mesh $N$ and a source image $X$, we obtain the expressive mesh with $M=\mathbf{D}_{\text{mesh}}(\mathbf{E}_{\text{exp}}(X),\mathbf{E}_{\text{id}}(N))$. For the other methods, we use their image encoders to obtain the expression coefficients $\boldsymbol{\Psi}$, and apply them directly to the neutral mesh with $M=N+\boldsymbol{\Psi}\boldsymbol{E}$. For FLAME-based models, this process is followed by the Linear Blend Skinning (LBS) function to account for jaw rotation. For experiments on ITW data neutrals are obtained with an off-the-shelf reconstruction method [12].
@@ 5.3 Results on MultiREX
We now show quantitative results on geometric expression capture, on the proposed MultiREX benchmark. We process videos from all subjects and camera views, comparing our model to DECA [16], EMICA [9, 10] and SMIRK [36]. The neutral mesh is manually selected (Sec. 5.1), and considered as a given for all approaches. Therefore MultiREX evaluation only accounts for geometric deformations due to the facial expressions. Additionally, we use ground truth bounding boxes, computed from projecting keypoint vertices to the image space. Results reported in bold are statistically significant based on a Wilcoxon signed rank test ($\rho<0.001$).
@@
We further analyze the viewpoint robustness of each method by reporting the average reconstruction errors for each viewing angle independently in Tab. 3. These results indicate that our model maintains performance better than competing methods as the viewing angle shifts from the frontal view. In contrast, the performance of EMICA and SMIRK decreases faster as the camera angle increases. This can be visualized in Fig. 6, where we show results for the same expression captured from five different views. Additional examples are provided in the supplementary material. Our model produces consistent results across views, particularly in side views when compared to competing methods. This is especially apparent in the mouth shape of the subject. This is explained by the fact that our method relies on a semantic space that represents facial expressions in their entirety, rather than relying solely on image space information where some regions may be occluded.
5.4 In-the-wild retargeting results
@@
In this section, we compare our method with recent state-of-the-art techniques, DECA [15], EMICA [10], SMIRK [36], 3DDFA-V3 [41], and TEASER [32] on the expression retargeting task from in-the-wild images. Since ground truth data for corresponding expressions across subjects on in-the-wild images is not possible to obtain, we compare methods qualitatively. We show several examples in Fig. 7, where the expressions from the images on the first column are transferred to the target identities of the second one.
@@
5.5 Ablations on the expression capture model
We conduct ablation studies to evaluate two key aspects of our approach: (i) the necessity of combining synthetic and real data, and (ii) the impact of the domain loss $L_{d}$. In Tab. 2, we see that training only on synthetic data leads to lower performance, indicating that mixing the synthetic and real domains is beneficial. Interestingly, still in Tab. 2, we see that the domain loss ($L_{d}$) slightly degrades performance on MultiREX. However, it is necessary for proper generalization to in-the-wild domain, as observed in Fig. 8. Results without $L_{d}$ exhibit limited motion and expressiveness, which is even more evident when viewing dynamic sequences in our supplementary video. $L_{d}$ introduces a trade-off between performance in a controlled lightstage environment and generalization to in-the-wild data. This trade-off is difficult to quantify, given that benchmarks for geometry accuracy like MultiREX do not exist in unconstrained settings.
Figure 8: Impact of the domain loss $L_{\text{d}}$ on in-the-wild capture.
@@ 6 Limitations
We focus on facial expression capture and consequently assume a known identity, therefore not estimating the neutral mesh. We acknowledge that the methods used for comparison in our experiments, while being the best candidates to our knowledge, were designed to address the broader task of reconstructing facial shape, expression and textures.
@@ 7 Conclusion
In this work, we present SEREP, an expression capture model that operates at the semantic level, enabling precise capture and retargeting of facial expressions. We start by learning a semantic expression model to represent expressions from unpaired data and directly in the 3D space. This model is leveraged to generate synthetic data with minimal manual labour, after sampling and retargeting known expression codes. Finally, we introduce a semi-supervised learning scheme that enables our expression capture model training, using the synthetic data and a domain adaptation strategy to generalize to unconstrained conditions.
Our experiments underscore the strengths of our expression representation, demonstrating its effectiveness in capturing expressions under various conditions, including challenging viewpoints and luminosity, while accounting for the identity facial geometry. Additionally, we introduce MultiREX, a new benchmark that evaluates expression geometry directly in 3D space. For future work, we aim to make our capture model temporally aware to improve video capture quality and plan to integrate camera position information for enhanced supervision.
@@ 8 Acknowledgements
This research was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) and MITACS through the Alliance-Mitacs program (Grant No. ALLRP 589317 - 23).
We would also like to thank authors of the Multiface dataset [47], as well as authors of DECA [16], EMICA [10], SMIRK [36], 3DDFAv3 [41], and TEASER [32] for releasing their code and advancing research in the field.
Appendices
Appendix A Implementation details
For training the semantic expression model, we use the following importance weight in eq. 1 of the main manuscript. $\lambda_{rec}=1.0$, $\lambda_{cycle}=1.0$, $\lambda_{delta}=0.01$, $\lambda_{edge}=10000$ and $\lambda_{eyes}=0.01$. Details for the model’s architecture are provided Tab. 4.
For the expression capture model, the following importance weight are used in eq. 2 of the main manuscript: $\lambda_{code}=10$, $\lambda_{lmks}=1$ and $\lambda_{domain}=0.005$. For the gradient reversal, we use a scale factor equal to 1. Details for the model’s architecture are provided Tab. 5.
Function Details
$\mathbf{E}_{\text{exp}}$ SpiralConv (x5) (3, 32) $\to$ (32, 32) $\to$ (32, 32) $\to$ (32, 64) $\to$ (64, 64) Linear: Linear(3392, 64)
$\mathbf{E}_{\text{id}}$ SpiralConv (x5) (3, 32) $\to$ (32, 32) $\to$ (32, 32) $\to$ (32, 64) $\to$ (64, 64) Linear: Linear(3392, 64)
$\mathbf{D}_{\text{mesh}}$ Linear: Linear(128, 3392) SpiralConv (x5) (64, 64) $\to$ (64, 32) $\to$ (32, 32) $\to$ (32, 32) $\to$ (32, 3)
Table 4: Model architecture for the semantic expression model.
Function Details
$\mathbf{E}_{\text{img}}$ ConvNeXt-B()
$\mathbf{H}_{\text{code}}$ ResBlock (x3): Linear(512, 512) $\to$ GELU $\to$ GroupNorm(32, 512) Linear: Linear(512, 64)
$\mathbf{H}_{\text{lmks}}$ Linear: Linear(512, 128)
$\mathbf{C}_{\text{d}}$ GradientReversal() Linear: Linear(512, 256) $\to$ GroupNorm(16, 256) $\to$ GELU ResBlock (x2): Linear(256, 256) $\to$ GELU $\to$ GroupNorm(16, 256) Linear: Linear(256, 1)
Table 5: Model architecture for the expression capture model.
Appendix B MultiREX additional details
@@
Appendix C Ablations on the semantic expression model
We ablate the eye closure loss $L_{\text{eyes}}$ and edge loss $L_{\text{edge}}$ in Fig. 10. Without $L_{\text{eyes}}$, eye closure is incomplete during blinks. The edge loss follows existing practice [2] and improves animations’ edge flow (removing jagged lines) to better support downstream animator needs. We will include these figures in the supplementary material.
Figure 10: Ablation on $L_{\text{eyes}}$ and $L_{\text{edge}}$.
@@ Appendix C Ablations on the semantic expression model
Appendix D More qualitative results
@@
D.1 Comparison on MultiREX benchmark
@@
D.2 Comparison on in-the-wild images
Fig. 15, we show more qualitative results for in-the-wild retargeting to other subjects and considering different source expressions. Our model is more faithful to the source expression in most scenarios while accounting for the face morphology and through its semantic expression model.
D.3 Challenging in-the-wild capture
We complete the evaluation of our expression capture model with challenging in-the-wild captures and some failure cases Fig. 16. Our model shows some robustness to lighting conditions, but can fail in cases of external occlusions, and motion blur. Motion blur is particularly problematic in video frames, resulting in jittery capture. The overall robustness of our method is in part dependent on the quality of the real landmarks used for training, obtained using an off-the-shelf detector.
Figure 16: Results on challenging conditions
Appendix E Real-time processing
On a 2080Ti GPU (batch size of 1), the expression encoder takes 20.6 ms on average, and the mesh decoder takes 5.5 ms, for a total of 26.1 ms (38.3 FPS). This does not include bounding box detection.
Appendix F Synthetic dataset
Fig. 17 shows samples from our synthetic dataset used to train our expression capture model. Rendering is done using Blender, with the Cycles renderer. We use randomly selected environment maps [^4] https://polyhaven.com. We use the same meshes and textures for teeth and eyes for all subjects. They are placed procedurally based on the vertex positions of the eyelids and jaw.